# AVA: Attentive VLM Agent for Mastering StarCraft II

### v3

@@ 4.3 Dynamic Role Assignment and Task Distribution
In multi-agent environments, particularly in complex dynamic scenarios like StarCraft II, efficient task allocation is fundamental to system performance optimization. Drawing inspiration from human societies’ division of labor and cooperation, we propose a novel VLM-Agent architecture that implements a unified approach to role assignment and task distribution, capable of adapting to rapidly evolving battlefield conditions while maintaining strategic coherence. Formally, let $\mathcal{N}=\{1,...,N\}$ denote the set of agents and $\mathcal{Z}=\{z_{1},...,z_{m}\}$ represent the set of possible roles. The role assignment function $\phi:\mathcal{N}\rightarrow\mathcal{Z}$ maps each agent to a specific role, while considering the state space $\mathcal{S}$ and action space $\mathcal{,T}$ of the environment. This mapping is dynamically optimized through a utility function $U(\phi,s)$ that evaluates the effectiveness of the role given the current state $s\in\mathcal{S}$.
Figure 3: Dynamic Role Assignment through VLM-Agent architecture.
The role assignment process incorporates a sophisticated reflection mechanism $\mathcal{M}$ that continuously evaluates performance through three key components:
1) Performance Monitoring: A real-time evaluation system that tracks key performance indicators (KPIs), including resource utilization efficiency, combat effectiveness, and strategic objective completion rates.
2) Historical Analysis: A temporal memory buffer that maintains a record of previous role assignments and their outcomes, enabling the system to learn from past experiences and adapt its strategies accordingly.
3) Adaptive Optimization: An iterative refinement process that adjusts role assignments based on both immediate feedback and historical performance patterns.
Our framework leverages a VLM through a sophisticated multimodal fusion function:
$z_{i}=\text{VLM}_{\text{role}}(I,T,C),$ (10)
where the model processes visual inputs $I$, textual prompts $T$, and contextual information $C$ to generate contextual representations. The system employs an iterative prompting mechanism with self-reflection, allowing the model to:
  • Generate comprehensive tactical assessments incorporating battlefield state, unit compositions, and resource distributions.
  • Evaluate potential outcomes through chain-of-thought reasoning to identify optimal strategies.
  • Refine decisions through structured prompt templates that encode tactical principles and maintain strategic consistency.
@@
Through this sophisticated prompt engineering framework, our VLM-Agent architecture achieves robust battlefield performance while maintaining interpretable decision-making processes. The system dynamically adapts to battlefield conditions by iteratively refining its prompts based on unit capabilities, environmental contexts, and emerging tactical situations. This approach effectively bridges the gap between abstract state representations and human-like strategic reasoning, while maintaining computational efficiency through structured prompt optimization.
The effectiveness of our approach is further enhanced by incorporating uncertainty estimation and risk assessment mechanisms, allowing the system to make more nuanced decisions in ambiguous situations. This is particularly crucial in scenarios where incomplete information or rapidly changing conditions require robust and adaptable decision-making capabilities.
@@ 4.4 Complete Pipeline
We formalize our decision-making pipeline as an iterative process that maps POMDP observations to actions through a series of VLM-based transformations. At each timestep $t$, given an observation $o_{t}=(I_{t},T_{t},U_{t})$ from the AVACraft environment state $s_{t}\in\mathcal{S}$, our system generates actions through the Algorithm 1.
The system maintains a history buffer $H_{t}$ for states and actions, and processes each step at 2Hz, maximizing the trade-off between strategic depth and real-time responsiveness. The sparse reward structure $R(s_{t})$ guides the objective of achieving victory while eliminating enemy units.
@@
5 Experiment
Table 2: Performance evaluation of AVA across different StarCraft II scenarios, used Qwen-VL-Plus. Results show varying success rates across different unit compositions and tactical situations, with higher performance in simpler scenarios and challenges in more complex engagements.
Map Scenario Win Rate (%)
vlm_attention_1 75
pvz_task6_level3 25
3m_vlm_attention 75
2s3z_vlm_attention 25
2s_vs_1sc_vlm_attention 0
2m_vs_1z_vlm_attention 10
2c_vs_64zg_vlm_attention 0
2bc1prism_vs_8m_vlm_attention 0
3s_vs_3z_vlm_attention 10
8marine_1medvac_vs_2tank 0
6reaper_vs8zealot_vlm_attention 0
8marine_2tank_vs_zerglings_ banelings_vlm_attention 25
Average Win Rate 20.4
Table 3: Performance comparison of different VLM models on two representative StarCraft II scenarios. Results demonstrate the relative capabilities of various models in handling both complex multi-unit scenarios (vlm_attention_1) and simpler tactical engagements (3m_vlm_attention).
Model Win Rate (%)
vlm_attention_1 3m_vlm_attention
GPT-4-Turbo 87 79
GPT-4o 79 81
GPT-4o-mini 76 76
Qwen-VL-Plus 75 75
Table 4: Ablation study on vlm_attention_1 map by GPT-4-Turbo. Checkmarks indicate which components are enabled in each configuration. Results demonstrate the individual and combined effects of Role Assignment, AVA mechanism, and Retrieval-Augmented Generation (RAG) on system performance.
Role Assignment AVA RAG Win Rate (%)
87
71
65
70
24
50
26
20
5.1 Experimental Setup
We evaluate our AVA framework across 12 specialized StarCraft II micromanagement scenarios. These scenarios are carefully designed to test different aspects of tactical decision-making, ranging from basic unit control to complex multi-unit coordination with ability usage.
For the VLM models, we primarily use GPT-4-Turbo as our main model due to its superior performance, while also testing with GPT-4o, GPT-4o-mini, and Qwen-VL-Plus for comparative analysis. Our VLM agents are built on Camel [^1] https://github.com/camel-ai/camel (Li et al., 2023), which supports convenient multi-agent system design and training. The environment implementation is built on PySC2222https://github.com/google-deepmind/pysc2 (Vinyals et al., 2017), with the environment running at 2Hz to balance decision quality and real-time performance.
5.2 Component Analysis Results
Our ablation study investigates the contribution of each key component in our framework: AVA mechanism, RAG system, and Role Assignment module. The results demonstrate strong synergistic effects between these components, with the complete system achieving the highest performance (87% win rate on vlm_attention_1).
AVA proves to be the most crucial component, particularly in complex scenarios requiring sophisticated target selection and tactical positioning. When enabled alone, it improves win rates by 30-40% compared to the baseline, highlighting its effectiveness in strategic decision-making.
The RAG system provides essential domain knowledge support, resulting in more consistent and tactically sound decisions. Its integration improves win rates by 20-25%, especially in scenarios requiring specific matchup understanding. Role Assignment further enhances coordination between units, contributing a 15-20% performance improvement when combined with other components.
5.3 Model Performance Analysis
Our comparative analysis of different VLM models reveals important trade-offs between performance, speed, and cost. GPT-4-Turbo achieves the best performance with win rates up to 87% but faces significant latency and cost challenges in real-time applications. GPT-4o and GPT-4o-mini offer better efficiency while maintaining reasonable performance (76-81% win rates).
Qwen-VL-Plus, despite being the most cost-effective option, shows limitations in knowledge consistency and instruction following. However, its accessibility makes it suitable for large-scale experiments and prototyping.
5.4 Scenario-Specific Performance
Our system demonstrates varying effectiveness across different scenario types. In basic engagement scenarios like 3m_vlm_attention and simple multi-unit coordination (vlm_attention_1), the system achieves high win rates (75%). These scenarios primarily test fundamental tactical decision-making and basic unit control.
However, performance degrades significantly in scenarios requiring precise ability usage or complex movement control. Maps requiring sophisticated micro-management, such as 2s_vs_1sc_vlm_attention and 2c_vs_64zg_vlm_attention, result in notably lower win rates (0-10%). This performance gap reveals current limitations in handling high-precision tactical maneuvers and ability timing.
5.5 Limitations and Challenges
Through our experimental evaluation, we identified several key challenges:
Spatial Understanding
VLMs show limited comprehension of StarCraft II’s spatial dynamics, particularly in scenarios requiring sophisticated terrain utilization. This limitation is evident in maps like 2c_64zg_vlm_attention (Figure 8), where opportunities for high-ground advantage are often missed.
Real-time Control
The high-frequency decision-making requirements of micromanagement pose significant challenges. Models often struggle with temporal consistency, exhibiting issues like excessive target switching and suboptimal movement commands.
Knowledge Application
We observe inconsistencies in domain knowledge application, variable output formatting, and occasional unit misidentification. These issues are particularly pronounced in complex scenarios requiring precise ability usage and unit coordination.
These limitations suggest that while our framework shows promise in strategic decision-making, significant improvements are needed in handling scenarios requiring precise micro-management and real-time control.
5.6 Emergent Capabilities
Through our experiments, we observe several remarkable emergent capabilities that VLM-based agents demonstrate without explicit training:
Strategic Unit Prioritization
In complex scenarios, particularly in vlm_attention_1 (Figure 7), agents demonstrate sophisticated tactical understanding by consistently prioritizing high-threat units (e.g., focusing Ghosts before they can use critical abilities) and low-health targets to maximize damage efficiency. This prioritization emerges naturally from the VLM’s understanding of unit capabilities and battlefield dynamics.
Multi-Unit Synergy
The system exhibits emergent coordination patterns, particularly evident in the 8marine_2tank_vs_zerglings_banelings_vlm_attention scenario (Figure 5). Here, the agent autonomously develops a sophisticated combined-arms strategy:
  • Marines automatically form protective formations around Siege Tanks.
  • Tanks are positioned to maximize splash damage against Banelings while minimizing friendly fire.
  • Unit positioning dynamically adjusts based on enemy movement patterns.
Adaptive Tactical Responses: The agents demonstrate the ability to develop complex micro-management skills without explicit programming:
  • Focus Fire: In 3m_vlm_attention, agents consistently identify and eliminate low-health targets to reduce incoming damage.
  • Formation Control: As demonstrated in the 8marine_1medvac_vs_2tank (Figure 6) through intelligent unit splitting to minimize splash damage.
  • Area Control: In 2c_vs_64zg_vlm_attention (Figure 4), Colossi are automatically positioned to maximize their line damage potential.
  • Resource Management: Systematic optimization of unit health and energy across various scenarios.
Context-Aware Decision Making: The system demonstrates an understanding of broader tactical contexts beyond immediate unit interactions:
  • Automatically adjusts aggression levels based on unit composition advantages.
  • Considers terrain features in positioning decisions.
These emergent capabilities are particularly noteworthy as they arise from the VLM’s general understanding of tactical principles, rather than from explicit training on StarCraft II mechanics. This suggests that VLM-based approaches may offer a promising direction for developing more flexible and adaptable game AI systems that can reason about complex tactical situations in ways similar to human players.
6 Conclusion
AVA represents a significant advancement in creating more human-like StarCraft II agents. By aligning artificial agent perception and decision-making processes with human cognition, our framework provides a more natural and effective environment for developing and testing AI systems. Future work will focus on expanding the system’s capabilities and evaluating its effectiveness in more complex scenarios.
7 Impact Statement
Appendix A Map Details

### v5

@@ 4.3 Dynamic Role Assignment and Task Distribution
In multi-agent environments, efficient task allocation is fundamental to system performance optimization. Drawing inspiration from human societies’ division of labor and cooperation, we implement a unified approach to role assignment and task distribution that adapts to rapidly evolving battlefield conditions while maintaining strategic coherence. Formally, let $\mathcal{N}=\{1,...,N\}$ denote the set of agents and $\mathcal{Z}=\{z_{1},...,z_{m}\}$ represent the set of possible roles. The role assignment function $\phi:\mathcal{N}\rightarrow\mathcal{Z}$ maps each agent to a specific role, while considering the state space $\mathcal{S}$ and action space $\mathcal{A}$ of the environment. This mapping is dynamically optimized through a utility function $U(\phi,s)$ that evaluates the effectiveness of the role given the current state $s\in\mathcal{S}$.
Figure 3: Dynamic Role Assignment framework showing the evaluation and optimization process.
The role assignment process incorporates a sophisticated evaluation mechanism $\mathcal{M}$ that continuously assesses performance through three key components: (1) Performance Monitoring - a real-time evaluation system that tracks key performance indicators (KPIs), including resource utilization efficiency, combat effectiveness, and strategic objective completion rates; (2) Historical Analysis - a temporal memory buffer that maintains a record of previous role assignments and their outcomes, enabling the system to learn from past experiences and adapt its strategies accordingly; and (3) Adaptive Optimization - an iterative refinement process that adjusts role assignments based on both immediate feedback and historical performance patterns. Our framework leverages VLMs through a multimodal fusion function:$z_{i}=\text{VLM}_{\text{role}}(I,T,C)$, where the model processes visual inputs $I$, textual prompts $T$, and contextual information $C$ to generate contextual representations. The system employs an iterative analytical process with continuous evaluation. This process enables comprehensive tactical assessments incorporating battlefield state, unit compositions, and resource distributions; evaluation of potential outcomes through structured reasoning to identify optimal strategies; and refinement of decisions through parameterized prompt templates that encode tactical principles while maintaining strategic consistency.
@@
Through this sophisticated prompt engineering framework, our architecture achieves robust battlefield performance while maintaining interpretable decision-making processes. The system dynamically adapts to battlefield conditions by iteratively refining its prompts based on unit capabilities, environmental contexts, and emerging tactical situations. This approach effectively bridges the gap between abstract state representations and human-like strategic reasoning, while maintaining computational efficiency through structured prompt optimization. The effectiveness of our approach is further enhanced by incorporating uncertainty estimation and risk assessment mechanisms, allowing the system to make more nuanced decisions in ambiguous situations. This is particularly crucial in scenarios where incomplete information or rapidly changing conditions require robust and adaptable decision-making capabilities.
@@ 4.4 Complete Pipeline
We formalize our decision-making pipeline as an iterative process that maps POMDP observations to actions through a series of VLM-based transformations. At each timestep $t$, given an observation $o_{t}=(I_{t},T_{t},U_{t})$ from the AVACraft environment state $s_{t}\in\mathcal{S}$, our system generates actions through the Algorithm 1 in Appendix C. The system maintains a history buffer $H_{t}$ for states and actions, and processes each step, maximizing the trade-off between strategic depth and real-time responsiveness. The sparse reward structure $R(s_{t})$ guides the victory while eliminating enemy units.
5 Experiments
We evaluate the Attentive VLM Agent (AVA) in the AVACraft environment to assess its tactical decision-making capabilities in StarCraft II micromanagement scenarios. Our experiments focus on two primary objectives: (1) quantifying AVA’s performance across diverse scenarios and vision-language models (VLMs), and (2) validating its alignment with human cognitive processes through expert evaluations. The implementation leverages the Camel framework (https://github.com/camel-ai/camel(Li et al., 2023) for multi-agent system design and PySC2 (https://github.com/google-deepmind/pysc2(Vinyals et al., 2017) for environment simulation, operating at 2Hz frequency to balance strategic decision depth and real-time responsiveness.
5.1 Quantitative Performance
We evaluated AVA across 12 micromanagement scenarios in AVACraft, designed to probe various tactical skills, including unit control, multi-unit coordination, and ability usage. These scenarios range from simple engagements (e.g., 3m) to complex battles requiring precise maneuvers (e.g., 2c_vs_64zg). We benchmarked four state-of-the-art VLM models—GPT-4-Turbo, GPT-4o, GPT-4o-mini, and Qwen-VL-Plus—with GPT-4-Turbo serving as the primary model due to its superior tactical reasoning capabilities.
Table 2: Performance comparison of different VLM models on two representative scenarios.
Model Win Rate (%)
mixed_units 3m
GPT-4-Turbo 87 79
GPT-4o 79 81
GPT-4o-mini 76 76
Qwen-VL-Plus 75 75
Table 3: Ablation study with GPT-4-Turbo.
Role MPI RAG Win Rate (%)
87
- 71
- 65
- 70
- - 24
- - 50
- - 26
- - - 20
Table 6 reports win rates for AVA with Qwen-VL-Plus, demonstrating strong performance (75%) in scenarios focused on strategic target selection like mixed_units and 3m, but revealing challenges in complex scenarios (0–10%) requiring precise ability timing or terrain exploitation. To systematically compare model performance, we evaluated all four VLMs on two representative scenarios (Table 2). GPT-4-Turbo achieved the highest performance with win rates of 87% (mixed_units) and 79% (3m), closely followed by GPT-4o (79% and 81%). To rigorously assess model robustness and tactical capabilities, we conducted a comprehensive head-to-head evaluation on mixed_units, pitting seven state-of-the-art VLMs against each other over 20 matches per pairing (Table 4). We conducted a comprehensive ablation study on vlm_priority_1 using GPT-4-Turbo (Table 3) to quantify the contributions of AVA’s architectural components.
The complete system integrating all three components (Role Assignment, Multimodal Priority Inference (MPI), and RAG) achieved an 87% win rate, with the MPI mechanism providing the most substantial individual contribution (50% win rate in isolation compared to the 20% baseline). The RAG component contributed 20-25% performance improvement through domain knowledge integration, while the Role Assignment framework added 15-20% through enhanced coordination capabilities, demonstrating the complementary nature of these architectural elements.
Table 4: Win-loss ratios in head-to-head matches between seven vision-language models on the vlm_priority_1 scenario, with 20 matches per pairing. Model abbreviations: GPT-4o-mini (gpt-4o-mini-2024-07-18), GPT-4o (gpt-4o-2024-08-06), GPT-4-Turbo (gpt-4-turbo-2024-04-09), Qwen-VL (qwen-vl-max), QVQ-Max (qvq-max-2025-03-25), Gemini-Flash (gemini-2.0-flash), Gemini-Lite (gemini-2.0-flash-lite).
Model
GPT-4o-mini
GPT-4o
GPT-4-Turbo
Qwen-VL
QVQ-Max
Gemini-Flash
Gemini-Lite
GPT-4o-mini 7:13 6:14 11:9 9:11 6:14 8:12
GPT-4o 13:7 9:11 13:7 11:9 9:11 10:10
GPT-4-Turbo 14:6 11:9 14:6 11:9 8:12 10:10
Qwen-VL 9:11 7:13 6:14 7:13 8:12 8:12
QVQ-Max 11:9 9:11 9:11 13:7 9:11 10:10
Gemini-Flash 14:6 11:9 12:8 12:8 11:9 11:9
Gemini-Lite 12:8 10:10 10:10 12:8 10:10 9:11
5.2 Human Evaluation and Interpretability
To assess AVA’s alignment with human cognitive processes and tactical reasoning, we conducted a structured evaluation with seven participants representing diverse StarCraft II expertise levels: one professional player, two Master-level players, one Diamond-level player, one Platinum-level player, one Gold-level player, one novice, and one spectator. Participants compared AVA against traditional MARL agents (https://www.youtube.com/watch?v=MLdqyyPcv9U) across three key metrics, each rated on a 1–5 scale: Game Bug Exploitation (higher scores indicating less exploitation), Reasoning Coherence, and Human Similarity (detailed metric definitions provided in Appendix E).
Table 5: Human evaluation comparing AVA and MARL approaches across three assessment metrics (1–5 scale): Bug Ex. (Game Bug Exploitation, higher is better), Reason. (Reasoning Coherence), Hum. Sim. (Human Similarity). Scores are stratified by evaluator expertise level: Expert (Professional, two Grandmasters, n=3), Mid-tier (Diamond, Platinum, Gold, n=3), and Novice/Spectator (n=2).
MARL AVA
Expertise Bug Ex. Reason. Hum. Sim. Bug Ex. Reason. Hum. Sim.
Expert 1.3 2.0 1.0 5.0 4.3 4.7
Mid-tier 2.0 2.0 1.3 3.7 4.7 4.7
Novice/Spect. 2.5 2.5 3.0 4.0 3.5 4.0
Average 1.9 2.1 1.7 4.3 4.3 4.5
Table 6: Win rates of AVA with Qwen-VL-Plus across 12 micromanagement scenarios.
Map Scenario Win Rate (%)
mixed_units 75
pvz_ht 25
3m 75
2s3z 25
2s_vs_1sc 0
2m_vs_1z 10
2c_vs_64zg 0
8m_vs_2pc1wp 0
3s_vs_3z 10
8m1mv_vs_2st 0
6r_vs_8z 0
8m2st_vs_35zg4b 25
Average Win Rate 20.4
As shown in Table 5, AVA significantly outperforms traditional MARL approaches across all evaluation metrics, achieving mean scores of 4.3, 4.3, and 4.5 for Game Bug Exploitation, Reasoning Coherence, and Human Similarity, respectively, compared to MARL’s substantially lower scores of 1.9, 2.1, and 1.7. This performance gap was particularly pronounced among expert evaluators, who gave AVA near-perfect scores for bug exploitation (5.0) and human similarity (4.7). Unlike MARL agents, which frequently exploit environment quirks by fixating on specific map positions (particularly evident in scenarios like 3s5z_vs_4s6z and corridor), AVA demonstrates sophisticated tactical behaviors that align with human gameplay strategies.
Expert evaluators specifically highlighted AVA’s implementation of advanced tactical principles, including armor-type targeting (e.g., Immortals prioritizing heavily-armored units, Archons focusing on Marine swarms) and professional-level micromanagement techniques such as focus-firing high-threat units like Banelings with Siege Tanks and executing hit-and-run tactics with Marines. These behaviors closely resemble high-level human gameplay strategies observed in professional matches (https://www.youtube.com/watch?v=WMawc4JQ5Dw at 33:54).
AVA’s emergent capabilities significantly enhance its interpretability and alignment with human cognitive processes. In scenarios like vlm_priority_1 (Figure 10), AVA demonstrates sophisticated target prioritization, identifying high-threat units (e.g., Ghosts with energy for special abilities) and low-health targets to maximize tactical advantage. In more complex scenarios like 8marine_2tank_vs_zerglings_banelings_vlm_priority (Figures 79), the system autonomously develops protective Marine formations around Siege Tanks, optimizing splash damage effectiveness while minimizing friendly fire risk. These complex behaviors emerge naturally from the Multimodal Priority Inference mechanism without explicit training, making AVA’s decision processes intuitive and interpretable even to non-expert observers.
6 Conclusion
AVA represents a significant advancement in creating more human-like StarCraft II agents. By aligning artificial agent perception with human cognition through RGB inputs and natural language processing, our framework bridges the gap between abstract state representations and human gameplay experience. The Multimodal Priority Inference mechanism, knowledge-enhanced decision system, and dynamic role assignment together enable complex tactical behaviors without explicit training. Experimental results demonstrate that AVA can execute sophisticated maneuvers while maintaining human-like decision processes, which traditional MARL methods struggle to achieve.
Our approach opens promising directions for future work, including improved spatial reasoning in dense formations and scaling to full-game scenarios. Beyond StarCraft II, AVA’s principles of multimodal perception and structured reasoning have broader implications for human-aligned AI in complex decision-making domains.
Appendix A Impact Statement
Appendix B Limitations of Previous StarCraft II Environments
While SMAC and SMACv2 have advanced multi-agent reinforcement learning research, they have fundamental limitations for developing AI systems that can truly master StarCraft II’s complex decision-making challenges:
Simplified Unit Abilities and Interactions
SMAC significantly simplifies unit abilities, removing critical micro-management elements that define StarCraft II gameplay. For example, Marines and Marauders lack Stimpack abilities, Stalkers cannot Blink, and only Medivacs retain their Heal ability. This oversimplification eliminates the rich tactical depth of StarCraft II, where ability timing and targeting often determine battle outcomes. In competitive play, a Marine without Stimpack is essentially a different unit, and skilled micro-management of these abilities is central to high-level play.
Limited Unit Diversity and Compositions
Both SMAC and SMACv2 feature extremely limited unit diversity, with most scenarios containing only 2-3 unit types. This fails to capture StarCraft II’s emphasis on complementary unit compositions and counter strategies. For instance, the classic ”Marine-Marauder-Medivac” composition requires specific control patterns that balance front-line positioning, focus fire, and healing priorities—tactical considerations absent in simplified environments.
Overly Simple Enemy AI
The enemy AI in SMAC and SMACv2 follows a basic ”attack spawn point” strategy without any tactical depth. It neither repositions units strategically nor prioritizes targets intelligently, creating unrealistic combat scenarios. This simplistic behavior fails to challenge agents to develop the sophisticated positioning and targeting skills needed in actual StarCraft II gameplay, resulting in strategies that don’t transfer to real matches.
Abstract State Representations
SMAC and SMACv2 represent the game state as abstract vectors containing unit attributes, positions, and health values, completely divorced from the visual and spatial reasoning humans use when playing. This misalignment between AI and human perception fundamentally limits the ecological validity of behaviors learned in these environments.
Questionable Randomization in SMACv2
While SMACv2 introduces procedural generation and randomization of unit types and positions, these changes don’t necessarily reflect meaningful tactical variations in StarCraft II. Random army compositions often create unrealistic scenarios that wouldn’t occur in competitive play, where army composition follows strategic principles and tech progression. This randomization tests an agent’s ability to handle arbitrary unit combinations but fails to evaluate tactical proficiency in realistic combat scenarios.
Focus on MARL Rather Than StarCraft II Mastery
These environments were designed specifically to advance MARL algorithms rather than to develop systems that can master StarCraft II gameplay. Consequently, they prioritize properties beneficial for reinforcement learning (like simplified action spaces and reward structures) over faithful reproduction of the tactical challenges that make StarCraft II compelling.
Our AVACraft environment addresses these limitations by preserving the rich tactical depth of StarCraft II micro-management. We maintain full unit abilities, support diverse unit compositions, create realistic combat scenarios, and—most importantly—align AI perception with human gameplay experience through RGB visual inputs and natural language observations. This approach enables the development of agents that can execute sophisticated tactical maneuvers involving ability timing, positioning, and multi-unit coordination that more closely resemble human gameplay.
Appendix C Pseudocode
@@
Appendix D Map Details